# Modeling information spread across networks with communities using a multitype branching process framework

### v1

@@ III.3 Cascade size distribution
In this section, we calculate the overall size of outbreaks, i.e., the cascade size distribution, with our framework. The cascade size distribution is also referred to as the total progeny distribution of a seed within the literature on branching processes kimmel_branching_2002 . This cascade size distribution can simply be found by introducing a counter type into our multi-type branching process that tracks the total number of active nodes that have been created for each community. [^2] The authors found the idea of using a counter in this manner in Ref. haccou2005branching , where it was introduced as a means of calculating expected population size in a MTBP. To achieve this, first let $C_{1}$ be a random variable tracking the total number of active nodes in a cascade, such that $P[C_{1}(t)=k]=P[\sum_{i}^{t}N_{1}(i)=k]$ and $C_{2}(t)$ is defined similarly as $P[C_{2}(t)=k]=P[\sum_{i}^{t}N_{2}(i)=k]$. The probability generation function that maps the joint probability distribution of variables $N_{1},N_{2},C_{1}$ and $C_{2}$ will be
@@
Here dummy variables $c_{1}$ and $c_{2}$ will track the cumulative number of active nodes in communities 1 and 2, respectively. The PGF $G_{\boldsymbol{C}(t)}$ is effectively an extension of $G_{N(t)}$, which includes two new counter types that track the number of active nodes in each community. Just as we had a recursive relationship $G_{N(t)}$ given by Eqn. (9), we have a recursive relationship for $G_{\boldsymbol{C}(t)}$ given by
$G_{\boldsymbol{C}(t)}(s_{1},s_{2},c_{1},c_{2})=G_{\boldsymbol{C}(t-1)}(G_{\boldsymbol{X^{(1)}}}(\overbrace{s_{1}c_{1}}^{\mathclap{\text{When a type-1 node is created a type-1 counter is created also.}}},s_{2}c_{2}),G_{\boldsymbol{X^{(2)}}}(s_{1}c_{1},s_{2}c_{2}),\underbrace{c_{1},c_{2}}_{\mathclap{\text{Number of counters persist from $t-1$ to $t$.}}})$ (26)
Here the counters for the number of nodes in community 1 or 2 are tracked by the dummy variables $c_{1}$ and $c_{2}$. Then each counter type always produce exactly one copy of themselves between generations, thus accounting for the total cascade size. The initial condition is then given as $G_{\boldsymbol{C}(0)}(s_{1},s_{2},c_{1},c_{2})=s_{1}c_{1}$, where we have one active node in community one ($s_{1}$) and a counter for that node ($c_{1}$).
As we are interested in the total cascade size, we substitute $c_{1}=c$ and $c_{2}=c$ in $G_{\boldsymbol{C}(t)}(1,1,c,c)$. We then marginalize $s_{1}$ and $s_{2}$ in Eqn. (25) to calculate the actual probabilities $P[C_{1}(t)+C_{2}(t)=n]$, which is a univariate PGF. The probability distribution can then be estimated from the univariate PGF using the inverse fast Fourier transform to obtain the cascade size distribution as discussed in App B. Moreover, in the same manner, functions $G_{C(t)}(1,1,c,1)$ and $G_{C(t)}(1,1,1,c)$ can yield the probability distribution for cascade sizes in community one and community two, respectively. As Fig. 3b) illustrates, our framework accurately estimates the expected cascade sizes not only for the entire network but also for each community individually.
@@ IV Extending the pgf framework to a general network with communities
In this section, we extend the analysis presented in Sec. III to a broader class of networks with community structure, where their degree distributions do not necessarily follow the Stochastic Block Model, i.e., the in- and between community degree distribution are not Poisson. While the networks constituting communities can be general, they are required to have a local tree-like structure and be well-described via the configuration model. In principle, it is possible to relax these assumptions, but this is left to future work.
Incorporating networks with a general degree requires careful consideration of the origin of node activation — whether from within or outside its own community. This in turn, will lead to the use of excess offspring distributions. Below, we explain how to incorporate this into our framework. As in the previous section, the first step is to derive the offspring distributions for each type of node we track. Let us start by calculating the offspring distribution for the seed node in community 1. For our framework to work we need to know (or be able to extract) the in-community degree of each node, $D^{(1)}_{1}$ and $D^{(2)}_{2}$, and the between-community degree of each node, $D^{(1)}_{2}$ and $D^{(2)}_{1}$. We then construct the PGFs corresponding degree distributions, e.g.,
@@
Now, let us consider what happens when the activations spread beyond the seed generation. According to ICM, a node cannot be activated again once it has been activated in a previous time step. If we have an active node, who was activated by a neighbour in the previous generation, the number of inactive (susceptible) nodes that can then go on to be activated differs from the seed generation. To account for this, we need to use the appropriate excess degree distribution.
@@
[…]
@@
To track each of these cases for community 1 we introduce two types, each representing where a node in community 1 was activated from. The same logic holds for community 2, therefore, we now must keep track of four distinct types of offspring produced, which leads us, as illustrated in Fig. 4, to a 4-type branching process. The corresponding PGFs will use dummy variables $s_{1}^{in}$ to track active nodes in community 1 that were activated through a link inside community 1, and $s_{1}^{out}$ for activations through links from community 2. Likewise, the offspring in community 2 obtained from travelling the edge inside and outside community 2 will be tracked by $s_{2}^{in}$ and $s_{2}^{out}$, respectively. Each of these types will have their own offspring distributions and probability-generating functions, which are shown in Tab. 1.
@@
$G_{\boldsymbol{X_{in}^{(1)}}}(s_{1}^{in},s_{2}^{out})$ $=\overbrace{G_{\widetilde{X}_{1}^{(1)}}(s_{1}^{in})}^{\mathclap{\text{Note use of excess offspring PGF due to }}\atop\mathclap{\text{parent coming from community 1.}}}\underbrace{G_{X_{2}^{(1)}}(s_{2}^{out})}_{\mathclap{\text{Offspring PGF due to parent coming }}\atop\mathclap{\text{ from community 1.}}}.$ (30)
@@ IV.1 Application of pgf framework to a log-normal network with community structure
To construct the required PGF we first need the in- and between-community degree distributions. To extract these, we simulate two networks of $5000$ nodes, each with a log-normal distribution internally with $\mu_{1}=0.85$ and $\sigma_{1}=0.4$ ($\mu_{2}=1.8$ and $\sigma_{2}=0.4$) for community 1 (and community 2, respectively). As the log-normal is a continuous distribution, the generate numbers are real-valued no integer valued, therefore, we apply the ceiling function to these sampled values. This yields an internal degree structure where the average degree of nodes in community 1 is approximately 3, and the average degree of nodes in community 2 is approximately 7. Figure 5a) shows the degree distribution for our example network. With these distributions, we can formulate the required PGFs $G_{X_{1}^{(1)}}$ and $G_{X_{2}^{(2)}}$, and then construct the excess offspring distributions $G_{\widetilde{X}_{1}^{(1)}}$ and $G_{\widetilde{X}_{2}^{(2)}}$.
@@ IV.2 The community seeding effect on cascades sizes
This section will focus on how the community structure, where an infection is seeded, changes the observed cascade size distribution. Figure 6 shows how the cascade size distribution changes for a range of infection ($\rho$) parameters when seeded in different communities, leaving the cross-community connection probability, $p_{e}$, fixed at $0.02$. We see that if we seed the branching process in community 1, which has a lower average degree, the probability of producing larger cascades is lower than if we seed initially in community 2 (which has a larger average degree). We can also note that our framework accurately captures this effect. See App. D for a wider parameter sweep and for how these values change when we also vary the between-community connection probability $p_{e}$.
Figure 6: Cascade size distribution changes for a range of infection ($\rho$) parameters when seeded in different communities, leaving the cross-community connection probability, $p_{e}$, fixed at $0.02$. Each community has $5000$ nodes, totalling $10000$ nodes in the network. The internal connections in community 1 and community 2 are those stated in Fig. 5. Lines represent theory curves, and the points represent the results of $N_{sim}=10^{6}$ Monte Carlo simulations of the ICM model on the network. Note that the middle dot-dashed line is for the corresponding simple branching process (which does not take into account community structure).
Next, we would like to isolate the effect of the community structure from the degree structure. To achieve this we compare the cascade dynamics on the network with communities to the cascades on the network with the same degree distribution but without community structure. To construct such a network, we take the network with community structure and perform a degree-preserving randomisation. This removes the community structure but preserves the degree structure as required. We then simulate the ICM on this network to compare it to our network with community structure.
@@
$D_{T}=\frac{1}{2}\overbrace{\left(D^{(1)}_{1}+D^{(1)}_{2}\right)}^{\mathclap{\text{The total degree of nodes in community 1.}}}+\frac{1}{2}\underbrace{\left(D^{(2)}_{1}+D^{(2)}_{2}\right).}_{\mathclap{\text{The total degree of nodes community 2.}}}$ (33)
@@
where $G_{\mathcal{N}_{s}(0)}(s)=s$. The PGF $G_{\mathcal{N}_{s}(1)}(s)$ provides us with a way of constructing a Simple Branching Process (SBP) that has the same overall degree distribution as our multi-type branching process but without community structure. From this model, we can derive cascade size distributions and plot them in Fig. 6 alongside the cascades on the corresponding network with communities. From Fig. 6, we see that the simple branching process model for a network without communities predicts a very different cascades size distribution than the model with communities. Thus the information diffusion on the community-based networks cannot be captured accurately with the Simple Branching Process.
@@
Figure 7: Cascade size distribution where $\rho$ is fixed to $0.12$. Solid and dashed lines represent $p_{e}=0.02$, where cross-community spread is possible, and dot-dashed lines represent $p_{e}=0$, where spread is confined to one community. a) Information spread behaviour when seeding in community 1 (which has an average degree of $\approx 3$). b) Information spread behaviour when seeding in community 2 (which has an average degree of $\approx 7$). As before, each community has $5000$ nodes, totalling $10,000$ nodes in the network. The internal connections in community 1 and community 2 use those stated in Fig. 5. Points represent the results of $N_{sim}=10^{6}$ Monte Carlo simulations of the ICM model on the network.
@@
One notable effect takes place when we seed a cascade in community 1 and we have a large enough probability of infection, $\rho=0.12$, the chance of observing larger cascades changes and displays quantitatively similar behaviour to those cascades seeded in community 2. We can speculate that this is due to activation probability being large enough that, if we observe a cascade of a certain size, it is likely to have spread to community 2, where the process takes advantage of the larger average degree within that community.
@@
We can easily isolate the effect of communities by removing the inter-community links, setting $p_{e}=0$ to verify our interpretations. This allows the activations to spread within either community 1 or 2 only; see Fig. 7. We can note from Fig. 7 that it is indeed due to the community structure that community 1 sees a change in its cascade size distribution; however, if we seed activations in community 2, we see no such discernible benefit for community 2 being a connection to community 1, i.e., there is little change in its cascade size distribution when it is connected to the other community.
@@
In this section, we have shown that our multi-type branching process model for information spread between communities can accurately model the spread of information between communities with a general network structure. This captures not only the effect of the community structure, where we have different internal connection structures within the communities, but also the effect of seeding infection in one community over the other. In the following section we provide some conclusions of our work and future directions.
@@ V Conclusion
In this paper, we studied the information spreading process on a broad class of networks with community structures. We showed how, by using limited information about the network, specifically the degree distribution within and between the communities, we can calculate standard statistical characteristics of cascade dynamics. For this purpose, we developed a probability generating function theory to study how information spreads between and within communities. The methods developed in this study add to the theoretical tools for studying information diffusion on community-based networks or multiple connected networks.
Our framework allowed us to calculate important quantities of information diffusion, such as extinction probability, hazard function, and cascade size distribution, proving to be a simple and effective tool. Not only did this tool enable us to accurately estimate the properties of dynamics on the entire network, but it also allowed us to extend these quantities to community-specific versions of characteristics of interest. This allowed us, for instance, to isolate the probability of extinction in a specific community. Additionally, we derive new quantities, specific to the information spread on a network with communities, such as the introduction and reintroduction probabilities. The introduction probability estimates (with high accuracy) the probability that the infection spreads to a new community, while the reintroduction represents the probability that a contagion, having ceased spreading in a specific community, might be reintroduced to that community at a later stage.
Finally, we analysed how information spreads across and within communities in a particular case in more details. We focused on a network that can be used to represent social networks, which was a graph with heavy-tailed (log-normal) distributed degree and sparsely distributed inter-community edges. We demonstrated how the initial seeding location of an infection affects the observed cascade size distribution. Furthermore, we compared the cascade dynamics on a network with communities to the dynamics on a network with the same degree distribution but no communities present. We showed that the presence of communities notably change the cascade dynamics on a heavy-tailed network and that our framework accurately captures the changes.
@@
To further investigate the effect of communities on information spread, it may be interesting to study cases involving communities of varying sizes and determine whether the size of a community influences cascade dynamics, as it reasonably would. Another intriguing scenario involves introducing different types of nodes with varying probabilities of infection. This addition will not significantly complicate our analysis but will increase the number of random variables we need to track within our framework.
To enhance the applicability of our method described here, a comparison to real world data is a natural extension. Previous work pena2023finding tracked information diffusion between and within communities for two conversation networks on Twitter (now X). The authors showed that users tend to communicate more frequently to other users in the same community than to users in the opposite community. But to what extent this is explained by the network structure and diffusion processes is an interesting and open question. The methodology introduced in this paper opens the door to the application to empirically observed cascades, to determined to what extent the network structure can explain the observed information diffusion process.

### v2

@@ III.3 Cascade size distribution
In this section, we calculate the overall size of outbreaks, i.e., the cascade size distribution, with our framework. The cascade size distribution is also referred to as the total progeny distribution of a seed within the literature on branching processes kimmel_branching_2002. This cascade size distribution can simply be found by introducing a counter type into our multi-type branching process that tracks the total number of active nodes that have been created for each community. The authors found the idea of using a counter in this manner in Ref. haccou2005branching, where it was introduced as a means of calculating the expected population size in a MTBP. We adapt it here for cascade distribution calculations.
To track the cascade size distribution we introduce $C_{1}$, a random variable tracking the total number of active nodes in a cascade from community one, such that $P[C_{1}(t)=k]=P[\sum_{i}^{t}N_{1}(i)=k]$ and, similarly, $C_{2}(t)$ is defined similarly as $P[C_{2}(t)=k]=P[\sum_{i}^{t}N_{2}(i)=k]$. Both $C_{1}(t)$ and $C_{2}(t)$ are tracked using new counter-types mapped to the dummy variables $c_{1}$ and $c_{2}$ within a PGF. These counter-type function as follows: every time a node becomes infected in community 1, a counter type, $c_{1}$, is created. Similarly, for every newly infected node in community 2, a counter type, $c_{2}$, is created. All counter-types are then forced to persist between generations. This translates to a probability generation function that maps the joint probability distribution of variables $N_{1},N_{2},C_{1}$ and $C_{2}$ which is given by
@@
The PGF $G_{\boldsymbol{C}(t)}$ is effectively an extension of $G_{N(t)}$, which includes the two new counter types that track the number of active nodes in each community. Just as we had a recursive relationship $G_{N(t)}$ given by Eqn. (9), we have a recursive relationship for $G_{\boldsymbol{C}(t)}$ given by
[…]
Note that each counter type $s_{1}$ and $s_{2}$ always produce exactly one copy of themselves between generations, thus accounting for the total cascade size. The initial condition is then given as $G_{\boldsymbol{C}(0)}(s_{1},s_{2},c_{1},c_{2})=s_{1}c_{1}$, where we have one active node in community one ($s_{1}$) and a counter for that node ($c_{1}$).
As we are interested in the total cascade size, we substitute $c_{1}=c$ and $c_{2}=c$ in $G_{\boldsymbol{C}(t)}(1,1,c,c)$. We then marginalize $s_{1}$ and $s_{2}$ in Eqn. (25) to calculate the actual probabilities $P[C_{1}(t)+C_{2}(t)=n]$, which is a univariate PGF. The probability distribution can then be estimated from the univariate PGF using the inverse fast Fourier transform to obtain the cascade size distribution as discussed in App. B. Moreover, in the same manner, functions $G_{C(t)}(1,1,c,1)$ and $G_{C(t)}(1,1,1,c)$ can yield the probability distribution for cascade sizes in community one and community two, respectively. As Fig. 3b) illustrates, our framework accurately estimates the expected cascade sizes not only for the entire network but also for each community individually.
@@ IV Extending the pgf framework to a general network with communities
In this section, we extend the analysis presented in Sec. III to a broader class of networks with community structure, where their degree distributions do not necessarily follow the Stochastic Block Model, i.e., the in- and between community degree distributions are not Poisson. While the networks constituting communities can be general, they are required to have a local tree-like structure and be well-described via the configuration model. In principle, it is possible to relax these assumptions, but this is left to future work.
Incorporating networks with a general degree requires careful consideration of the origin of node activation — whether from within or outside its own community. This in turn, will lead to the use of excess offspring distributions. Below, we explain how to incorporate this into our framework. As in the previous section, the first step is to derive the offspring distributions for each type of node we track. Let us start by calculating the offspring distribution for the seed node in community 1. For our framework to work we need to know (or be able to extract) the in-community degree of each node, $D^{(1)}_{1}$ and $D^{(2)}_{2}$, and the between-community degree of each node, $D^{(1)}_{2}$ and $D^{(2)}_{1}$. We then construct the PGFs corresponding to degree distributions, e.g.,
@@
Now, let us consider what happens when the activations spread beyond the seed generation. According to ICM, a node cannot be activated again once it has been activated in a previous time step. If we have an active node, who was activated by a neighbor in the previous generation, the number of inactive (susceptible) nodes that can then go on to be activated differs from the seed generation. To account for this, we need to use the appropriate excess degree distribution.
@@
    If it was activated from inside its own community, then to correctly track activations from it, we need to use the excess offspring distribution for community 1, with associated random variable $\widetilde{X}_{1}^{(1)}$, (as we have used up one possible link that the activations could travel down), and the offspring distribution for between-communities, with associated random variable $X_{2}^{(1)}$, (as we have not used up any edges between communities from which activations could traverse).
@@
To track each of these cases for community 1 we introduce two types, each representing where a node in community 1 was activated from. The same logic holds for community 2, therefore, we now must keep track of four distinct types of offspring produced, which leads us, as illustrated in Fig. 4, to a 4-type branching process. The corresponding PGFs will use dummy variables $s_{1}^{in}$ to track active nodes in community 1 that were activated through a link inside community 1, and $s_{1}^{out}$ for activations through links from community 2. Likewise, the offspring in community 2 obtained from traveling the edge inside and outside community 2 will be tracked by $s_{2}^{in}$ and $s_{2}^{out}$, respectively. Each of these types will have their own offspring distributions and probability-generating functions, which are shown in Tab. 1.
@@
[…]
@@ IV.1 Application of pgf framework to a log-normal network with community structure
To construct the required PGF we first need the in- and between-community degree distributions. To extract these, we simulate two networks of 5 0005\,000 nodes, each with a log-normal distribution internally with $\mu_{1}=0.85$ and $\sigma_{1}=0.4$ ($\mu_{2}=1.8$ and $\sigma_{2}=0.4$) for community 1 (and community 2, respectively). As the log-normal is a continuous distribution, the generate numbers are real-valued no integer valued, therefore, we apply the ceiling function to these sampled values. This yields an internal degree structure where the average degree of nodes in community 1 is approximately 3, and the average degree of nodes in community 2 is approximately 7. Figure 5a) shows the degree distribution for our example network. With these distributions, we can formulate the required PGFs $G_{X_{1}^{(1)}}$ and $G_{X_{2}^{(2)}}$, and then construct the excess offspring distributions $G_{\widetilde{X}_{1}^{(1)}}$ and $G_{\widetilde{X}_{2}^{(2)}}$.
@@ IV.2 The community seeding effect on cascades sizes
This section focuses on how the community structure, in which an infection is seeded, affects the observed cascade size distribution. Figure 6 shows how the cascade size distribution changes for a range of infection ($\rho$) parameters when seeded in different communities, leaving the cross-community connection probability, $p_{e}$, fixed at $0.02$. We see that if we seed the branching process in community 1, which has a lower average degree, the probability of producing larger cascades is lower than if we seed initially in community 2 (which has a larger average degree). We can also note that our framework accurately captures this effect. See App. D for a wider parameter sweep and for how these values change when we also vary the between-community connection probability $p_{e}$.
Figure 6: Cascades across the entire network when seeded in different communities for a range of infection ($\rho$) parameters. Here the cross-community connection probability, $p_{e}$ is fixed at $0.02$. Each community has 5 0005\,000 nodes, totaling 10 00010\,000 nodes in the network. The internal connections in community 1 and community 2 are those stated in Fig. 5. Lines represent theory curves, and the points represent the results of $N_{sim}=10^{6}$ Monte Carlo simulations of the ICM model on the network. Note that the middle pink line is for the corresponding simple branching process (which does not take into account community structure).
Next, we would like to isolate the effect of the community structure from the degree structure. To achieve this we compare the cascade dynamics on the network with communities to the cascades on the network with the same degree distribution but without community structure. To construct such a network, we take the network with community structure and perform a degree-preserving randomization. This removes the community structure but preserves the degree structure as required. We then simulate the ICM on this network to compare it to our network with community structure.
@@
[…]
@@
where $G_{\mathcal{N}_{s}(0)}(s)=s$. The PGF $G_{\mathcal{N}_{s}(1)}(s)$ provides us with a way of constructing a Simple Branching Process (SBP) that has the same overall degree distribution as our multi-type branching process but without community structure. From this model, we can derive cascade size distributions and plot them in Fig. 6 alongside the cascades on the corresponding network with communities. From Fig. 6, we see that the simple branching process model for a network without communities predicts a very different cascade size distribution from the model with communities. Thus,diffusion on community-based networks cannot be captured accurately with the Simple Branching Process. The use of our MTBP framework is necessary.
@@
Figure 7: Cascades across the entire network when $\rho$ is fixed to $0.12$. Purple and blue lines (MTBP) represent $p_{e}=0.02$, where cross-community spread is possible, and pink lines (SBP) represent $p_{e}=0$, where spread is confined to one community. a) Diffusion behavior when seeding in community 1 (which has an average degree of $\approx 3$). b) Diffusion behavior when seeding in community 2 (which has an average degree of $\approx 7$). Here the blue and pink lines coincide. As before, each community has 5 0005\,000 nodes, totaling 10 00010\,000 nodes in the network. The internal connections in community 1 and community 2 use those stated in Fig. 5. Points represent the results of $N_{sim}=10^{6}$ Monte Carlo simulations of the ICM model on the network.
@@
One notable effect takes place when we seed a cascade in community 1 and we have a large enough probability of infection, $\rho=0.12$, the chance of observing larger cascades changes and displays quantitatively similar behavior to those cascades seeded in community 2 (i.e., the shape deviates and appears similar to cascade size distribution when cascades are seeded in community 2). We can speculate that this is due to activation probability being large enough that, if we observe a cascade of a certain size, it is likely to have spread to community 2, where the process takes advantage of the larger average degree within that community. In contrast, we see no similar effect in community 2.
@@
We can easily isolate the effect of communities by removing the inter-community links, setting $p_{e}=0$ to verify our interpretations. This allows the activations to spread within either community 1 or 2 only. Fig. 7a) shows the cascade size distribution when information is seeded within community 1 and either allowed to spread between communities or not, while Fig. 7b) shows the cascade size distribution when information is seed within community 2 and either allowed to spread between communities or not. We can note from Fig. 7a) that it is indeed due to the community structure that community 1 sees a change in its cascade size distribution, where the inter community links provide a means of linking to the higher, on average, degree nodes in community 2. However, if we seed activations in community 2, we see no such discernible benefit for community 2 being a connection to community 1, where the inter community links provide a means of linking to the lower, on average, degree nodes in community 1, which does not have a qualitative effect on the cascade size distributions shape. The interested reader can see the effect of communities whose average degree is closer to each others in App. F.
@@
In this section, we have shown that our multi-type branching process model for diffusion between communities can accurately model the spread of entities between communities with a general network structure. This captures not only the effect of the community structure, where we have different internal connection structures within the communities, but also the effect of seeding infection in one community over the other. In the following section we provide some conclusions of our work and future directions.
@@ V Conclusion
In this paper, we studied the spreading process on a broad class of networks with community structures. We showed how, by using limited information about the network, specifically the degree distribution within and between the communities, we can calculate standard statistical characteristics of cascade dynamics. For this purpose, we developed a probability generating function theory to study how the diffusion process spreads between and within communities. The methods developed in this study add to the theoretical tools for studying diffusion on community-based networks or multiple connected networks.
Our framework allowed us to calculate important quantities of the diffusion process, such as extinction probability, hazard function, and cascade size distribution, proving to be a simple and effective tool. Not only did this tool enable us to accurately estimate the properties of dynamics on the entire network, but it also allowed us to extend these quantities to community-specific versions of characteristics of interest. This allowed us, for instance, to isolate the probability of extinction in a specific community. Additionally, we derived new quantities, specific to diffusion on a network with communities, such as the introduction and reintroduction probabilities. The introduction probability estimates (with high accuracy) the probability that the infection spreads to a new community, while the reintroduction represents the probability that a contagion, having ceased spreading in a specific community, might be reintroduced to that community at a later stage.
Finally, we analyzed how the diffusion process spreads across and within communities in a particular case in more details. We focused on a network that can be used to represent social networks, which was a graph with heavy-tailed (log-normal) distributed degree and sparsely distributed inter-community edges. We demonstrated how the initial seeding location of an infection affects the observed cascade size distribution. Furthermore, we compared the cascade dynamics on a network with communities to the dynamics on a network with the same degree distribution but no communities present. We showed that the presence of communities notably change the cascade dynamics on a heavy-tailed network and that our framework accurately captures the changes.
In this work, we have used the Independent Cascade Model (ICM) as the diffusion process. Although very useful for biological contagions, it may not describe how information spreads on a real social network. In gleeson_branching_2021, it was noted that the empirical cascade size distribution of social spreads was not captured well by the ICM. However, simple extensions of the ICM performed very well. For instance, the Limited Attention Model lerman2016information; weng2012competition scales the probability of a node spreading a piece of information by the number of links they have. This scaling attempts to capture the cognitive load of users who may see many pieces of content from many users that they are in contact with. Such a model would be a natural extension to this work, which would better reflect real-world spread dynamics than the ICM and is an interesting future direction.
@@
To further investigate the effect of communities on propagation dynamics, it may be interesting to study cases involving communities of varying sizes and determine whether the size of a community influences cascade dynamics, as it reasonably would. Another intriguing scenario involves introducing different types of nodes with varying probabilities of infection. This addition will not significantly complicate our analysis but will increase the number of random variables we need to track within our framework.
To enhance the applicability of our method described here, a comparison to real world data is a natural extension. Previous work pena2023finding tracked information diffusion between and within communities for two conversation networks on Twitter (now X). The authors showed that users tend to communicate more frequently to other users in the same community than to users in the opposite community. But to what extent this is explained by the network structure and diffusion processes is an interesting and open question. The methodology introduced in this paper opens the door to the application to empirically observed cascades, to determine to what extent the network structure can explain the observed information diffusion process.